Rana thinks we can do better than 18 W heat load on the vacuum can. The insulation seems to have good cover all the way around but I want to see objectively where heat is leaking out.
The vacuum can heaters have been hooked up to a 30 V ~0.5A supply (P=V^2/R=30^2/50=18W). This will be left overnight so that the assembly can come up to temperature. I will have a look tomorrow with a thermal imaging camera to see if heat leakage points can be spotted.
The RTD 100 Ohm thermal sensor that was hooked up to the 966EN-4006 acromag card had been moved to monitor the air temperature within the tent. The LIGO temperature sensor interface board (see chain of posts starting at PSL:1730) was reconnected to power, the pinouts are tabulated below. One of the AD590 sensors (LEMO#6) was routed to the 8th input of the Acromag ADC a new channel was created in the data base (C3:PSL-VACTEMP_MON1) to monitor the vacuum can temperature. I'm not entirely sure if Frank selected good metal film resistors for this circuit (see PSL:843), he was only using this for diagnostic monitoring. Might need to switch to the J5 later as this one appears to at a calibrated value that makes sense when I tested it.
A frame builder would be good right now to log this data.
Pin outs of the DSub-9 connectors on the Temp sensor board: